# Cluster Categorification of Rank 2 Webs

### v1

@@ 6. Rank 2 webs as cluster variables
The cluster structure on $\rm Conf_{4}\mathcal{A}$ is quasi-isomorphic to the one on the double Bruhat cell $G^{w_{0},w_{0}}$ (here $w_{0}$ is the longest word in the Weyl group $S_{k}$ of the Lie group ${\rm SL}_{k}$). The cluster structure on $G^{w_{0},w_{0}}$ was studied extensively by Berenstein, Fomin and Zelevinsky [BFZ05] and this quasi-isomorphism is a folklore; it is implicit in Fock-Goncharov [FG06], see also [LL23].
There they show that cluster variables can be written down using the combinatorics of double reduced words. We consider the Weyl group $S_{k}\times S_{k}$, where we generate the first factor by simple reflections $11,2,\dots,k-$ 1 and we generate the second factor by simple reflections $\overline{1},\overline{2},\dots,\overline{k-1}$. One chooses a double reduced word for $(w_{0},w_{0})$, which is a word in the alphabet $1,2,\dots,k-1,\overline{1},\overline{2},\dots,\overline{k-1}$. Corresponding to any such word is a cluster:
@@
Let $i=a+c-k$. Note that $b=(k-a)+(k-c)$. Now, let us choose a double reduced word which contains as a partial word some permutation $(u,v)$ such that
@@
Then the function $f_{u\omega_{i},v\omega_{i}}$ depends on $a$-form $v_{1}\wedge\dots\wedge v_{a}$ from $F_{1}$, the $(k-a)$-form $v_{a+1}\wedge\dots\wedge v_{k}$ from $F_{2}$, the $(k-c)$-form $v_{k+1}\wedge\dots\wedge v_{2k-c}$ from $F_{3}$, and the $c$-form $v_{2k-c+1}\wedge\dots\wedge v_{2k}$ from $F_{4}$:
@@ 7. Cactus sequence and Main Result
(see [FL19]) Let $C$ be a cluster on $\rm Conf_{3}\mathcal{A}$. Then there exists a cluster $C^{\prime}$ on $\rm Gr(k,3k-3)$ such that $\{\phi^{*}(f)|f\in C\}\subset C^{\prime}$. Moreover the quiver for $C^{\prime}$ restricts to the quiver for $C$. Thus we can view the cluster structure on $\rm Conf_{3}\mathcal{A}$ as “living inside” the cluster structure on $\rm Gr(k,3k-3)$.
@@
Now let us define the functions appearing in the cactus sequence. Let $a,b,c,d\leq k$, $a>d$, $a+b+c+d=2k$. Moreover, assume that $a-d,b,c$ satisfy the triangle inequality. For any such $a,b,c,d$ we define the function ${\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}$ on $\rm Conf_{3}\mathcal{A}$. It is determined by its pullback to the Grassmannian $\rm Gr(k,3(k-1))$, $\phi^{*}({\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}})$. Because we will be primarily working with functions on the Grassmannian, let us simply call this pullback function ${\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}$ as well.
@@
    The function ${\begin{pmatrix}a&\\ &c\\ b&\end{pmatrix}}$ for $a+b+c=k$ and $a,b,c\leq k-1$ is the Plücker coordinate involving $21,2,\dots,a,k,k+1,\dots k+b-1,2k-1,2k,\dots,2k+c-$ 2.
@@
Figure 9. Rank $1$ webs.
@@
The following webs correspond to the functions $\bigl{(}\begin{smallmatrix}4,2&\\ &2\\ 2&\end{smallmatrix}\bigr{)}$, $\bigl{(}\begin{smallmatrix}4&\\ &4\\ 2&\end{smallmatrix}\bigr{)}$, $\bigl{(}\begin{smallmatrix}3,1&\\ &3\\ 3&\end{smallmatrix}\bigr{)}$ in $\rm Gr(5,12)$. Notice that, up to the leaves intersecting, these webs are trees, so the fact that they are cluster variables in $\rm Gr(5,12)$ confirms the philosophy of Fomin and Pylyavskyy [FP16, Conjecture 10.1]. We can see that these functions, via stretching, can live in many Grassmanians, and they will be cluster variables thanks to Theorem 6.6.
Figure 10. Rank $2$ webs.
@@
[…]
@@
The particulars of this mutation sequence, like the precise sequence of mutations and the clusters that occur along the way, are not so important for us. We will only use that each mutation which has the following form:
@@
[[Le19] Proposition 3.2]
@@
Note that in Proposition 7.4, all the the mutations involve rank $2$ functions. However, we can use the degeneracy to view rank $1$ functions (Plücker coordinates) as rank $2$ functions: Thus we start with the function ${\begin{pmatrix}d&\\ &c\\ b&\end{pmatrix}}={\begin{pmatrix}k,d&\\ &c\\ b&\end{pmatrix}}$ and mutate until we end up with the function ${\begin{pmatrix}k-d,0&\\ &c+d\\ b+d&\end{pmatrix}}$.
@@
Because the functions in this sequence are all of the ones in $\mathcal{W}$, up to stretching, it is enough to understand the modules attached to the functions in this sequence. Our strategy is to use the cactus mutation sequence to recursively verify that the module corresponding to $W\in\mathcal{W}$ is $\Psi(W)$ for all the functions ${\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}$.
@@
are the ones coming from the map $\Psi$.
From this, we would like to compute the module which categorifies ${\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}}$. In order to do this, we will show that $\Psi{\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}}$ satisfies the appropriate exact sequences:
Theorem 7.5.
@@
[…]
@@
    the previous maps are compatible and determin a map $U\hookrightarrow V\oplus W$.
@@
Thus once we know that the functions ${\begin{pmatrix}5,1&\\ &4\\ 4&\end{pmatrix}},\Psi{\begin{pmatrix}4,1&\\ &4\\ 5&\end{pmatrix}},\Psi{\begin{pmatrix}5&\\ &5\\ 4&\end{pmatrix}},\Psi{\begin{pmatrix}4,1&\\ &5\\ 4&\end{pmatrix}},\Psi{\begin{pmatrix}5&\\ &4\\ 5&\end{pmatrix}}$ all correspond to the modules given by applying $\Psi$, we see that this is also the case for ${\begin{pmatrix}4&\\ &5\\ 5&\end{pmatrix}}$.
@@
But by 5.4, we understand how stretching of functions corresponds to stretching of modules: if $\Psi{W}$ is the module corresponding to $W$, this holds for any web obtained from stretching $W$. Thus for all $W\in\mathcal{W}$, $\Psi(W)$ is the module corresponding to $W$.

### v2

@@ 6. Rank 2 webs as cluster variables
The cluster structure on $\rm Conf_{4}\mathcal{A}$ is quasi-isomorphic to the one on the double Bruhat cell $G^{w_{0},w_{0}}$ (here $w_{0}$ is the longest word in the Weyl group $S_{k}$ of the Lie group ${\rm SL}_{k}$). The cluster structure on $G^{w_{0},w_{0}}$ was studied extensively by Berenstein, Fomin and Zelevinsky [BFZ05] and this quasi-isomorphism is folklore; it is implicit in Fock-Goncharov [FG06], see also [LL23].
In [BFZ05], they show that cluster variables can be written down using the combinatorics of double reduced words. We consider the Weyl group $S_{k}\times S_{k}$, where we generate the first factor by simple reflections $1,2,\dots,k-1$ and we generate the second factor by simple reflections $\overline{1},\overline{2},\dots,\overline{k-1}$. One chooses a double reduced word for $(w_{0},w_{0})$, which is a word in the alphabet $1,2,\dots,k-1,\overline{1},\overline{2},\dots,\overline{k-1}$. Corresponding to any such word is a cluster:
@@
Let i=a+cki=a+c-k. Note that b=(ka)+(kc)b=(k-a)+(k-c). Now, let us choose a double reduced word which contains as a partial word some permutation $(u,v)$ such that
@@
Then the function $f_{u\omega_{i},v\omega_{i}}$ depends on $a$-form $v_{1}\wedge\dots\wedge v_{a}$ from $F_{1}$, the (ka)(k-a)-form $v_{a+1}\wedge\dots\wedge v_{k}$ from $F_{2}$, the (kc)(k-c)-form $v_{k+1}\wedge\dots\wedge v_{2k-c}$ from $F_{3}$, and the $c$-form $v_{2k-c+1}\wedge\dots\wedge v_{2k}$ from $F_{4}$:
@@
Therefore, the modules $\Psi(W)$ for $W\in\mathcal{W}$ are all reachable. It is an open question whether all rigid indecomposable modules are reachable via mutation. Thus Corollary 6.7 suggests the possibility that all rigid indecomposable modules are reachable by mutation.
@@ 7. Cactus sequence and Main Result
(see [FL19]) Let $C$ be a cluster on $\rm Conf_{3}\mathcal{A}$. Then there exists a cluster $C^{\prime}$ on $\rm Gr(k,3k-3)$ such that $\{\phi^{*}(f)|f\in C\}\subset C^{\prime}$. Moreover the quiver for $C^{\prime}$ restricts to the quiver for $C$. Thus we can view the cluster structure on $\rm Conf_{3}\mathcal{A}$ is a cluster substructure of the cluster structure on $\rm Gr(k,3k-3)$.
@@
Now let us define the functions appearing in the cactus sequence. Let $a,b,c,d\leq k$, $a>d$, $a+b+c+d=2k$. Moreover, assume that ad,b,ca-d,b,c satisfy the triangle inequality. For any such $a,b,c,d$ we define the function $\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}$} on $\rm Conf_{3}\mathcal{A}$. It is determined by its pullback to the Grassmannian $\rm Gr(k,3(k-1))$, $\phi^{*}({\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}})$. Because we will be primarily working with functions on the Grassmannian, let us simply call this pullback function $\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}$} as well.
@@
    The function $\begin{pmatrix}a&\\ &c\\ b&\end{pmatrix}$} for $a+b+c=k$ and $a,b,c\leq k-1$ is the Plücker coordinate involving $1,2,\dots,a,k,k+1,\dots k+b-1,2k-1,2k,\dots,2k+c-2$.
@@
Figure 9. Rank 11 webs.
@@
The following webs correspond to the functions $\bigl(\begin{smallmatrix}4,2&\\ &2\\ 2&\end{smallmatrix}\bigr)$, $\bigl(\begin{smallmatrix}4&\\ &4\\ 2&\end{smallmatrix}\bigr)$, $\bigl(\begin{smallmatrix}3,1&\\ &3\\ 3&\end{smallmatrix}\bigr)$ in $\rm Gr(5,12)$. Notice that, up to the leaves intersecting, these webs are trees, so the fact that they are cluster variables in $\rm Gr(5,12)$ confirms the philosophy of Fomin and Pylyavskyy [FP16, Conjecture 10.1]. We can see that these functions, via stretching, can live in many Grassmannians, and they will be cluster variables thanks to Theorem 6.6.
Figure 10. Rank 22 webs.
@@
$\begin{pmatrix}4&\\ &\\ 1&\end{pmatrix}$$\begin{pmatrix}3&\\ &\\ 2&\end{pmatrix}$$\begin{pmatrix}2&\\ &\\ 3&\end{pmatrix}$$\begin{pmatrix}1&\\ &\\ 4&\end{pmatrix}$$\begin{pmatrix}4&\\ &1\\ &\end{pmatrix}$$\begin{pmatrix}3&\\ &1\\ 1&\end{pmatrix}$$\begin{pmatrix}2&\\ &1\\ 2&\end{pmatrix}$$\begin{pmatrix}1&\\ &1\\ 3&\end{pmatrix}$$\begin{pmatrix}&\\ &1\\ 4&\end{pmatrix}$$\begin{pmatrix}3&\\ &2\\ &\end{pmatrix}$$\begin{pmatrix}2&\\ &2\\ 1&\end{pmatrix}$$\begin{pmatrix}1&\\ &2\\ 2&\end{pmatrix}$$\begin{pmatrix}&\\ &2\\ 3&\end{pmatrix}$$\begin{pmatrix}2&\\ &3\\ &\end{pmatrix}$$\begin{pmatrix}1&\\ &3\\ 1&\end{pmatrix}$$\begin{pmatrix}&\\ &3\\ 2&\end{pmatrix}$$\begin{pmatrix}1&\\ &4\\ &\end{pmatrix}$$\begin{pmatrix}&\\ &4\\ 1&\end{pmatrix}$}
@@
The particulars of this mutation sequence, like the precise sequence of mutations and the clusters that occur along the way, are not so important for us. What is important for us is that each mutation has the following form:
@@
[LE19, Proposition 3.2]
Every mutation in the cactus sequence transforms cluster variables
@@
Note that in Proposition 7.4, all the mutations involve rank $2$ functions. However, we can use the degeneracy to view rank $1$ functions (Plücker coordinates) as rank $2$ functions: Thus we start with the function ${\begin{pmatrix}d&\\ &c\\ b&\end{pmatrix}}={\begin{pmatrix}k,d&\\ &c\\ b&\end{pmatrix}}$ and mutate until we end up with the function $\begin{pmatrix}k-d,0&\\ &c+d\\ b+d&\end{pmatrix}$}.
@@
Because the functions in this sequence are all of the ones in $\mathcal{W}$, up to stretching, by Lemma 5.8 it is enough to understand the modules attached to the functions in this sequence. Our strategy is to use the cactus mutation sequence to recursively verify that the module corresponding to $W\in\mathcal{W}$ is $\Psi(W)$ for all the functions $\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}$}.
@@
are the ones coming from the map $\Psi$. From this, we would like to compute the module which categorifies $\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}$}.
In order to do this, we will first show that $\Psi{\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}}$ sits in the appropriate exact sequences:
Proposition 7.5.
@@
$0\rightarrow\Psi{\begin{pmatrix}5,1&\\ &4\\ 4&\end{pmatrix}}\xrightarrow{f}\Psi{\begin{pmatrix}4,1&\\ &5\\ 4&\end{pmatrix}}\oplus\Psi{\begin{pmatrix}5&\\ &4\\ 5&\end{pmatrix}}\xrightarrow{g}\Psi{\begin{pmatrix}4&\\ &5\\ 5&\end{pmatrix}}\rightarrow 0$
@@
    the previous maps are compatible and determine a map $U\hookrightarrow V\oplus W$.
@@
Thus once we know that the functions ${\begin{pmatrix}5,1&\\ &4\\ 4&\end{pmatrix}},{\begin{pmatrix}4,1&\\ &4\\ 5&\end{pmatrix}},{\begin{pmatrix}5&\\ &5\\ 4&\end{pmatrix}},{\begin{pmatrix}4,1&\\ &5\\ 4&\end{pmatrix}},{\begin{pmatrix}5&\\ &4\\ 5&\end{pmatrix}}$ all correspond to the modules given by applying $\Psi$, we see that this is also the case for $\begin{pmatrix}4&\\ &5\\ 5&\end{pmatrix}$}.
@@
Proposition 7.6.
For the clusters appearing in the cactus sequence, the irreducible maps come in four types (up to degeneration):
  • (i)
    $\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}\rightarrow\Psi{\begin{pmatrix}a-1,d&\\ &c+1\\ b&\end{pmatrix}}$
  • (ii)
    $\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}\rightarrow\Psi{\begin{pmatrix}a,d-1&\\ &c\\ b+1&\end{pmatrix}}$
  • (iii)
    $\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}\rightarrow\Psi{\begin{pmatrix}a,d+1&\\ &c-1\\ b&\end{pmatrix}}$
  • (iv)
    $\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}\rightarrow\Psi{\begin{pmatrix}a+1,d&\\ &c\\ b-1&\end{pmatrix}}$
These maps are given by the maps in Figure 12 up to stretching.
Figure 12. Four different irreducible maps in the cactus sequence.
Proof.
These maps can be tracked through the cactus sequence inductively. As explained in [LE19, Proposition 3.2], there are four types of mutations in the sequence including degenerations. For each of these types of mutations, one should check if the irreducible maps are as stated in the proposition before and after the mutation. This is a finite check.
@@
The mutation sequences in Proposition 7.5 are non-split exact sequences giving minimal left (respectively right) $\rm add$}-$(T/M)$-approximations of $M=\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}$ for the appropriate cluster-tilting object $T$.
In other words, they allow us to inductively conclude that if $\Psi{\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}}$ categorifies $\begin{pmatrix}a,d&\\ &c\\ b&\end{pmatrix}$}, then $\Psi{\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}}$ categorifies $\begin{pmatrix}a-1,d-1&\\ &c+1\\ b+1&\end{pmatrix}$}. From this we may conclude:
Theorem 7.8.
@@
But by Lemma 5.8, we understand how stretching of functions corresponds to stretching of modules: if $\Psi(W)$ is the module corresponding to $W$, this holds for any web obtained from stretching $W$. Thus for all $W\in\mathcal{W}$, $\Psi(W)$ is the module corresponding to $W$.